At the beginning of the year, Smith Company budgeted overhead of $129,600 as well as 13,500 direct labor hours. During the year, Job K456 was completed with the following information: direct materials cost, $2,750; direct labor cost, $5,355. The average wage for Smith Company employees is $17 per hour.

By the end of the year, 18,100 direct labor hours had actually been worked, and Smith incurred the following actual overhead costs for the year:

Equipment lease $ 6,800
Depreciation on building 19,340
Indirect labor 90,400
Utilities 14,560
Other overhead 41,400
Required:
1. Calculate the overhead rate for the year.
2. Calculate the total cost of Job K456.
3. Prepare the Dec. 31 journal entries to record actual overhead and to apply overhead to production for the year.
4. Is overhead overapplied or underapplied? By how much?
5. Assuming that the normal cost of goods sold for the year is $635,600, what is the adjusted cost of goods sold?
Answer #1

1) Overhead rate = estimated overhead/estimated hour = 129600/13500 = 9.60 per hour

2) Total cost of Job K456

Direct material 2750
Direct labor 5355
Overhead (5355/17*9.60) 3024
Total cost of Job 11129

3) Journal entry

Date account and explanation Debit Credit
Manufacturing overhead 172500
Lease payable 6800
Accumulated depreciation-Building 19340
Wages payable 90400
Utilities payable 14560
Account payable 41400
(To record actual overhead)
Work in process (18100*9.6) 173760
Manufacturing overhead 173760
(To record applied overhead)

4) Over applied overhead = 173760-172500 = 1260

5) Adjusted cost of goods sold = 635600-1260 = 634340